# IOVS4NeRF: Incremental Optimal View Selection for Large-Scale NeRFs

### v1

@@ Abstract
Urban-level three-dimensional reconstruction for modern applications demands high rendering fidelity while minimizing computational costs. The advent of Neural Radiance Fields (NeRF) has enhanced 3D reconstruction, yet it exhibits artifacts under multiple viewpoints. In this paper, we propose a new NeRF framework method to address these issues. Our method uses image content and pose data to iteratively plan the next best view. A crucial aspect of this method involves uncertainty estimation, guiding the selection of views with maximum information gain from a candidate set. This iterative process enhances rendering quality over time. Simultaneously, we introduce the Vonoroi diagram and threshold sampling together with flight classifier to boost the efficiency, while keep the original NeRF network intact. It can serve as a plug-in tool to assist in better rendering, outperforming baselines and similar prior works.
Index Terms:
Uncertainty Estimation, UAV, Neural Radiance Fields, Scene Reconstruction, View Selection
@@ I Introduction
The essence of image-based 3D reconstruction is to infer the geometry and structure of objects and scenes from single or multiple 2D images. This longstanding challenge with broad implications is fundamental to many applications from scene understanding to object recognition, including medical diagnosis, robot navigation, 3D modeling and industrial control. The task of 3D reconstruction has indeed matured, as evidenced by commercial products such as Agisoft and Pix4D, which are capable of delivering high-fidelity, large-scale 3D models.
There’s a growing trend to harness deep learning algorithms inspired by animal vision[1, 2]. Exploring the fusion of 3D reconstruction and deep learning has revealed new possibilities, yet significant challenges persist.For instance, Point clouds[3, 4], consists of points with 3D coordinates, potentially incorporating color and reflectance intensity information. However, it lacks surface connectivity and the reconstructed surfaces appearing less smooth.
In contrast to previous methods, Neural Radiance Fields (NeRF)[5] utilizes neural networks to represent the radiance field in a scene. NeRF employs a simple multi-layer perceptron (MLP) network, and volumetric rendering for novel view synthesis and image reconstruction. It utilizes a straightforward five-dimensional input comprising three-dimensional spatial coordinates ($x,y,z$) and two-dimensional ray directions ($\theta,\phi$) to ultimately produce color and voxel density for images. Despite its efficacy, NeRF’s implicit neural network results in resource-intensive computations[6] and demands substantial pose information for training[7, 8], leading to inefficiencies in large scenes.
@@
Figure 1: Novel view synthesis comparison. We propose a new NeRF framework, and the comparision on the ArtSci dataset across four different model strategies, revealing that our method outperforms others in terms of both processing time and reconstruction efficacy.
View Planning. Our view selection in this article is based on view planning[9]. It involves determining the optimal angles, positions, and camera or sensor settings for capturing relevant information. Various visual tasks necessitate distinct viewpoint planning algorithms, broadly categorized into search-based and synthesis-based methods. Search-based methods involve sampling numerous candidate views and selecting specific views under certain constraints, whereas synthesis methods significantly reduce computational costs but lack sufficient reliability and accuracy.
The Next-Best view (NBV) problem[10] is a specific subfield of view planning. In NBV, algorithms assess information gain or target visibility from different viewpoints, selecting the most valuable or informative view to advance the task. Requirements for search-based viewpoint planning include: (1) sampling a specific number of candidate views in the view space (2) simulating visual information for each candidate view and estimating the information gain (3) selecting the optimal viewpoint or a set of viewpoints.
@@
Information Gain. Our method quantifies selection criteria through information gain. Originally, information entropy is utilized to describe the uncertainty and randomness of information, representing the complexity of a random variable. Information gain is defined as the amount of information a feature can bring to a classification system; higher information gain indicates greater feature importance.
In the NeRF-NBV problem[11], uncertainty quantification is often expressed as a representation of uncertainty. Existing research frequently employs uncertainty quantification results as the expression of information gain. In deep learning, uncertainty estimation methods mainly include Bayesian estimation (as seen in S-Nerf[12]), MC Dropout[13], and model ensembles[14]. Model ensembles reduce the bias and variance of a single model by combining predictions from multiple models. Expanding on this, given the context of drone flight, we integrate rendering uncertainty and position uncertainty as novel standard reference. We treat both with equal importance to pinpoint the next best view that maximizes information gain.
IOVS4NeRF. In our NeRF method, tailored processing methods are applied based on distinct flight trajectories. Initially, an adaptive classifier is trained to discern whether a given aerial trajectory, captured through images, is planar or non-planar. Despite the majority of drone photography trajectories could be seen planar[15, 16], we enhance the classifier’s generalization capability. For planar trajectories, we compress and confine position uncertainty within Voronoi information radiation field. Conversely, for non-planar trajectories, a Voronoi clustering algorithm is employed to derive uncertainty from the distances between different viewpoints and a central point.
Given the challenges posed by the lengthy training time and insufficient preservation of detailed information inherent in position encoding and MLP implicit networks, we use Instant-NGP[17] as a tool to assist us better render our scene. In the context of volume rendering, it has been demonstrated[18, 6]that NeRF struggles to generalize effectively from a limited number of input views. When the scene observations are incomplete, the original NeRF framework tends to collapse to trivial solutions by predicting zero volume density for unobserved regions. To address this, we model the radiance color at each position in the scene as a beta distribution, aligning with its mathematical properties between 0 and 1. This modeling approach not only simplifies computations but also provides a better fit for large scenes, adding a crucial refinement to the overall framework method, as mentioned in Figure 1.
In summary, our contributions are as follows.
• We propose IOVS4NeRF, a novel NeRF approach for efficient 3D scene reconstruction in large-scale environments. Using a lightweight network architecture and enhanced volumetric rendering techniques, our method achieves significant improvements in both training and rendering efficiency.
• We use rendering uncertainty together with position uncertainty as the mixed criterion proxy for the total information gain of the underlying 3D geometry, and we present an uncertainty guided policy using NeRF representation for next-best-view selection.
• Our strategy, while maintaining reconstruction effectiveness, significantly reduces the time required for 3D reconstruction. Extensive experiments on real-world datasets have been performed to show the efficiency of the proposed method.
@@ II Related Work
Traditional Real Scene Reconstruction. Traditional 3D modeling data acquisition mainly includes manual modeling and oblique photogrammetry. Due to the labor-intensive nature and low efficiency of manual modeling tools such as 3DSMax and Sketch Up, oblique photogrammetry with drones equipped with five-lens cameras capturing image data from multiple angles to obtain complete and accurate texture data and position information has a slight advantage. However, its reconstruction still follows three steps. The first is Structure from Motion (SfM[19]), which analyzes feature points in image sequences, utilizes bundle adjustment parameters to obtain 3D structure and camera trajectories. This is followed by Multi-View Stereo(MVS[20]), which estimates depth maps for each viewpoint to generate a single-viewpoint cloud. Finally, surface reconstruction merges different viewpoint clouds and conducts surface reconstruction[21].
With the development of deep learning, scholars have integrated this method into 3D modeling. DeepVO[22] uses deep recursive CNN to directly infer poses from a series of RGB images, bypassing traditional visual odometry modules. BA-Net[23] incorporates the Bundle Adjustment (BA) optimization algorithm as a layer in a neural network. This is done to train a more efficient set of basis functions, simplifying the backend optimization process during reconstruction.
Uncertainty Estimation. In computer vision, there’s a growing recognition of the crucial role uncertainty estimation plays. It not only enhances the interpretability of neural network outputs but also mitigates the risks of critical errors in important tasks. This trend reflects a heightened focus on the reliability of deep learning models. A detailed analysis of uncertainty in predictions offers a more thorough understanding of model behavior in complex real-world scenarios, paving the way for more reliable solutions.
This development primarily comprises two components: Bayesian Neural Networks (BNN)[24] and Dropout Variational Inference[25]. BNN treats the weights of neural networks as parameters of probability distributions, introducing prior and posterior distributions to infer the uncertainty in model outputs. Dropout Variational Inference, on the other hand, utilizes dropout layers for multiple inferences, estimating uncertainty by sampling different subnetworks.
Furthermore, research explores the application of uncertainty estimation in the domain of novel view synthesis. NeRF-W[26] introduces uncertainty to model transient objects in a scene, focusing on differences between images rather than inherent noise in training data, as observed in ActiveNeRF[27]. Tri-MipRF[28] approximates cones with a multivariate Gaussian, constructing inscribed spheres for sample points. It projects these spheres orthogonally onto three planes, obtaining circular projections.
NoPe-NeRF[29] utilizes only color and depth uncertainty without considering camera pose. The novel depth loss proposed in this paper learns scale and shift parameters for inter-frame depth, equivalent to maintaining geometrically consistent scenes during learning. However, these complex training and rendering processes can be daunting. Thus, we propose rendering uncertainty, combining position uncertainty. Additionally, in the domain of unmanned aerial vehicles, we innovatively incorporate Voronoi diagram[30, 31] to quantify position uncertainty, significantly reducing training and rendering time.
Fast Training and Rendering in Large Scenes. City-level radiance fields, exemplified by approaches like CityNeRF[32], Mega-NeRF[33], and Grid-NeRF[34], represent significant advancements tailored specifically for urban environments. Urban Radiance Fields effectively leverage RGB images and LiDAR scan data for 3D reconstruction and novel view synthesis. Subsequent papers build upon this foundation, incorporating Deformable Neural Mesh Primitives (DNMP[35]) as a novel carrier. Initially, DNMP is generated from point clouds, and through rasterization, ray-DNMP intersection points are computed. Feature interpolation is employed, coupled with view-dependent embedding, to determine radiance values and opacity .
CityNeRF employs a progressive learning strategy, activating high-frequency channels in position encoding to preserve intricate details during training. On the other hand, BlockNeRF[36] trains individual blocks within scene subdivisions, enabling fine-tuning block by block. Mega-NeRF utilizes NeRF++[37] for vast, unbounded scenes and enhances background expressiveness. It also incorporates the appearance embedding method from NeRF-W to balance varying lighting conditions across different images within the same scene.
Grid-guided Neural Radiance Fields[34] propose a two-branch model, consisting of a pretraining Grid branch stage and a Joint Learning Grid-and-NERF branches stage. During the joint learning stage, grid branch loss and nerf branch loss are jointly optimized. The sampling points are positioned closer to object surfaces, resulting in impressive outcomes. These methods collectively propel advancements in understanding and rendering urban-level scenes, fostering applications within urban environments.
@@
Figure 3: Overview of our framework. Our pipeline consists of three parts: I. Based on the 5D information of the images, we input them into NeRF for training. We define the color at a certain position as a beta distribution and quantitatively obtain the mean and variance of color from the MLP, generating our rendering uncertainty. II. We extract the positional information of the images from the training set and input them into a classifier to determine their respective Voronoi positional uncertainty, resulting in hybrid uncertainty. III. Based on the hybrid information uncertainty, we select the images with the maximum gain from the candidate set and add them to the training set, iterating continuously.
@@ III Background
NeRF represents a scene as a continuous function $F_{\theta}$ that outputs both the emitted radiance value and volume density. Specifically, given a 3D position $\mathbf{x}=(x,y,z)$ in the scene and a viewing direction vector $\mathbf{d}=(\theta,\phi)$, a multi-layer perceptron model is used to generate the corresponding volume density $\sigma$ and color $\mathbf{c}=(r,g,b)$ as follows:
@@
, where $\gamma_{x}(\cdot)$ and $\gamma_{d}(\cdot)$ are the positional encoding functions, and $f$ represents the intermediate feature that is independent of the viewing direction $\mathbf{d}$. An interesting observation is that the radiance color is influenced solely by its own 3D coordinates and the viewing direction, rendering it independent of other locations.
@@
$\hat{C}(\mathbf{r})$ $=\sum_{i=1}^{N_{s}}\alpha_{i}c(\mathbf{r}(t_{i}))$ (4)
$\alpha_{i}$ $=\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right)\left(1-\exp(-\sigma_{i}\delta_{i})\right)$
,where $\delta_{i}=t_{i+1}-t_{i}$ is the distance between adjacent samples, and $N_{s}$ denotes the number of samples.Based on this approach, NeRF optimizes the continuous function $F_{\theta}$ by minimizing the squared reconstruction errors between the ground truth RGB images $\{I_{i=1}^{N}\}$ and the rendered pixel colors.
@@
, where $\mathbf{r_{i}}$ is a sampled ray, and $C(\mathbf{r_{i}}),\hat{C}^{c}(\mathbf{r_{i}}),\hat{C}^{f}(\mathbf{r_{i}})$ correspond to the ground truth, coarse model prediction, and fine model prediction respectively.
@@ IV IOVS4NeRF
In this paper, we focus on optimal view selection for large-scale NeRFs. In this section, we present the IOVS4NeRF framework in section. IV-A followed by detailed descriptions of the proposed hybrid-uncertainty estimation (section. IV-B) and implementation (section. IV-C).
@@ IV-A Framework
1) Initialization. As shown in Fig. 2, we randomly select a fixed proportion of images (we experimentally recommend 15%) from the input dataset to perform NeRF training for initialization.
2) Uncertainty estimation. Note that the IOVS4NeRF selects the best view based on the information gain metric, named hybrid-uncertainty, with two main terms: 1) rendering uncertainty and 2) positional uncertainty:
For rendering uncertainty, we input 5D coordinates (x,d) of the remaining photo set into a modified NeRF network with threshold sampling, which then outputs both color c and volume density variance. Subsequently, we calculate the rendering uncertainty of the image by integrating both c and variance into a beta distribution.
For positional uncertainty, we input 3D coordinates x of the remaining photo set, and a classifier determines whether the trajectory is planar or non-planar. We then estimate the position information using Voronoi diagrams to obtain the positional uncertainty.
Then, we normalize and sum the two uncertainty to calculate the hybrid uncertainty, as shown in the following formula:
${\psi^{2}}(I)=Norm(\sigma_{pos}^{2}(I))+Norm(\sigma_{rgb}^{2}(I))$ (6)
, I means the ground truth RGB images and the function $Norm(\cdot)$ representing normalization. The formula $\sigma_{pos}^{2}(I)$ and $\sigma_{rgb}^{2}(I)$ can been seen at Eq. (17) and Eq. (7).
We select the image with the highest hybrid uncertainty and continually add it to the training set. This process is repeated until a specific reconstruction effect is achieved or a preset limit on the number of selected images is reached.
We partition the hybrid uncertainty into rendering uncertainty and position uncertainty. Kendall and Gal[38] identified two types of uncertainty relevant to computer vision—aleatoric uncertainty and epistemic uncertainty, and our uncertainty belongs to the latter type, which can be reduced with more data.
$\sigma_{rgb}^{2}(I)=\sum\limits_{i=1}^{{N_{r}}}{\frac{{||C({{\bf{r}}_{i}})-\bar{C}({{\bf{r}}_{i}})||_{2}^{2}}}{{2{{\bar{\beta}}^{2}}({{\bf{r}}_{i}})}}+\frac{{\log{{\bar{\beta}}^{2}}({{\bf{r}}_{i}})}}{2}}$ (7)
, where $N_{r}$ means the total number of the rays in photo I.
@@
IV-B Hybrid-Uncertainty Estimation
$\bullet$ Rendering Uncertainty.
Focused on scenarios with limited training data, if the scene observation is incomplete, the original NeRF framework often collapses into fragmented results by predicting zero volume density for unobserved regions. As a remedy, we propose modeling the color rgb value at each position in the scene as a beta distribution[39, 40] rather than a single value. The predicted variance can reflect the uncertainty at a particular position. This approach allows the model to provide larger variances in unobserved regions, thus maintaining greater information gain for completing renderings and reducing the overall model’s epistemic uncertainty.
Moreover, by constraining color values between 0 and 1, in line with the mathematical characteristics of the beta distribution, its conjugate prior inherently includes the partition function in the Bayesian formula, thereby avoiding complex global integrals in Bayesian inference. Thus, it can be quantitatively calculated and scaled in large-scale scenes.
Specifically, we define the color at a certain position as a beta distribution with parameters representing the mean $c(\textbf{r}({t_{i}}))$ and variance ${{\beta}^{2}}(\textbf{r}(t_{i}))$. Following Bayesian neural networks, we use the model output as the mean and add an additional branch outside the MLP network of NeRF to establish the variance model:
${\beta^{2}}({\bf{r}}({t_{i}}))=-P({\alpha_{i}})\log P({\alpha_{i}})$ (8)
, where ${\alpha_{i}}$ is the same as Eq. (4), and $P(\cdot)$ represents proportion of ${\alpha_{i}}$in the $\sum{{\alpha_{i}}}$ along the ray.
During the rendering process, similar volume rendering methods can be used to handle new neural radiation fields with uncertainty. The design paradigm in the NeRF framework provides two key prerequisites. The first is that the volume density at a specific location is only influenced by its own 3D coordinates, not affected by the viewing direction $\mathbf{d}$, which makes the distributions at different positions independent of each other. The second is that volume rendering can be approximated as a linear combination of sampled points along rays. Based on these conditions, if we represent the beta distribution at the position $\textbf{r}(t_{i})$ then, according to the conjugacy of the beta distribution, the rendering values along this ray naturally follow a beta distribution as:
$\hat{C}({\textbf{r}})\sim Beta(\bar{C}(\textbf{r}),{{\bar{\beta}}^{2}}(\textbf{r}))$ (9)
, where
$\bar{C}({\textbf{r}})=\sum\limits_{i=1}^{{N_{s}}}{{\alpha_{i}}c({\textbf{r}}({t_{i}}))}$ (10)
${\bar{\beta}^{2}}({\textbf{r}})=\sum\limits_{i=1}^{{N_{s}}}{\alpha_{i}^{2}{\beta^{2}}({\textbf{r}}({t_{i}}))})$ (11)

### v2

@@ Abstract
Neural Radiance Fields (NeRF) have recently demonstrated significant efficiency in the reconstruction of three-dimensional scenes and the synthesis of novel perspectives from a limited set of two-dimensional images. However, large-scale reconstruction using NeRF requires a substantial amount of aerial imagery for training, making it impractical in resource-constrained environments. This paper introduces an innovative incremental optimal view selection framework, IOVS4NeRF, designed to model a 3D scene within a restricted input budget. Specifically, our approach involves adding the existing training set with newly acquired samples, guided by a computed novel hybrid uncertainty of candidate views, which integrates rendering uncertainty and positional uncertainty. By selecting views that offer the highest information gain, the quality of novel view synthesis can be enhanced with minimal additional resources. Comprehensive experiments substantiate the efficiency of our model in realistic scenes, outperforming baselines and similar prior works, particularly under conditions of sparse training data.
@@ I Introduction
Neural Radiance Fields(NeRFs) [1] has significantly enables the perception of 3D scenes using 2D data[2]. Compared to previous deep learning approaches [3, 4], such as those based on point clouds [5, 6], NeRF offers higher reconstruction efficiency and remarkable adaptability in different research realms. It employs an implicit MLP network to better perceive the physical world[7] and has found applications in various fields, including robotics and 3D modeling.
The original NeRF encodes scene information into network parameters and uses differentiable volume rendering, which requires substantial computational resources [8], especially for large-scale scenes, leading to inefficient reconstruction. To address these issues, previous works [9, 10, 11, 12, 13] have utilized combinations of hash functions, spherical harmonics, and both explicit and implicit scene representations to enable faster training of NeRF. Some works have also expanded NeRF reconstruction to focus on large-scale scenes [14, 15, 16, 17]. These methods mainly involve decomposing a large scene into several smaller scenes, with each small scene independently trained by a NeRF sub-module. While this approach can achieve high-fidelity reconstruction and significantly increase reconstruction speed compared to the original NeRF, it is based on the premise of using extensive resources. However, in resource-constrained situations where only a limited number of images can be used as input, the reconstruction results tend to have severe artifacts and lower visual fidelity.
@@
Figure 1: Novel View Synthesis Comparison. We propose a new NeRF framework, and the comparison on the ArtSci dataset across four different model strategies reveals that our method outperforms others in terms of both processing time and reconstruction efficiency.
Inspired by the Next-Best View (NBV) problem [18], we reduce the number of images in the dataset to minimize computational resource consumption. We aim for each image in the limited input to contain as much information as possible, meaning the dataset is composed of images with the highest information gain to maximize scene understanding. Selecting the next best viewpoint to maximize information gain typically involves using uncertainty estimation for quantification. To address the problem of training NeRF in resource-constrained scenarios, some recent works have employed uncertainty-guided NBV selection[19, 20, 21, 22]. These methods significantly modify the internal network, resulting in low reusability. Other works[23] do not make any changes to NeRF itself, offering strong generalization capabilities, but mostly consider image uncertainty only during the rendering process. Although these methods have achieved good results, they still face challenges with artifacts and efficiency when applied to large-scale scenes.
Based on the above series of work, we model the radiance color at each sampling point in the scene as a Gaussian distribution to solve the above uncertainty references, which is used to assess rendering issues caused by unseen regions, including content outside the scene observed. For large-scale scene problems, inspired by oblique photography[24], we divide UAV flight paths into planar and non-planar trajectories[25, 26]. For these two different trajectories, we define a positional uncertainty to evaluate the positional gain provided by different 3D poses in real-world scenarios. We integrate rendering uncertainty and positional uncertainty as a novel standard metric hybrid uncertainty. We treat both with equal importance to determine the next best view that maximizes information gain.
Additionally, by selecting lightweight NeRF models, we further accelerate the reconstruction speed from the perspective of model structure. In this paper, we use Instant-NGP[9] as a tool to help us better reconstruct our scenes. Ultimately, we aim to achieve high-quality and high-efficiency NeRF reconstruction in large scenes with fewer computational resources. In summary, our contributions are as follows:
@@
• We propose a NeRF framework for incremental optimal view selection, aiming at addressing the limitations of NeRFs in large-scale scenes with restrained computational cost.
• We define a novel estimation of uncertainty, which is comprised of rendering uncertainty and positional uncertainty, to compute the information gain of each candidate view.
• We extensively evaluate our method on real-world datasets, showing it significantly reduces time while maintaining reconstruction effectiveness compared to previous baselines.
@@ II Related Work
II-A Neural Radiance Fields
NeRF [1] is a high-quality scene reconstruction technique that combines 3D reconstruction with neural rendering technology. Its main task is novel view synthesis, which involves capturing a series of images from several known viewpoints and synthesizing images from unseen viewpoints. This method encodes a 3D scene as a radiance field represented by MLP. The network is used to learn the volumetric density and color of the 3D scene, optimizing the continuous volumetric scene and modeling it with high fidelity. Due to the encoding of scene information in the network parameters and the computation of differentiable volume rendering, NeRF’s training and rendering processes consume substantial computational resources and time. As a result, many variations of the original NeRF have emerged for efficient reconstruction [9, 10, 11, 12, 13]. Among these methods, Instant-NGP uses multiresolution hash encoding, which is adaptive and efficient. It not only accelerates training and conserves resources but also has good adaptability to scenes of different scales. Therefore, we chose to base our research on it in this paper.
II-B NeRFs for Large-Scale Scene
Due to the advantages of high-quality reconstruction offered by NeRF, various NeRF methods for large-scale scene reconstruction have emerged [14, 15, 16, 17]. Given the limited capacity of the original NeRF, which is unable to handle the vast data volumes in large-scale scenes, researchers have primarily adopted the idea of decomposing large scenes into several smaller ones, each with a corresponding NeRF module. Notable examples of NeRF methods that utilize scene partitioning include Mega-NeRF proposed by Turki et al. [15] and Block-NeRF proposed by Tancik et al. [16]. Mega-NeRF decomposes large-scale scenes captured by drones into multiple different spatial sub-scenes and uses a geometric clustering algorithm to divide the input images into different NeRF submodules that can be trained in parallel. Block-NeRF, on the other hand, divides large street scenes captured by vehicles into different small scene blocks, training each block in segments and fine-tuning them sequentially.
Datasets used for large-scale scene reconstruction often contain numerous images, which consume significant computational resources. Under resource constraints, reducing the number of dataset images can help decrease the computational load. However, applying these methods with fewer images directly often results in severe artifacts and lower visual fidelity. The challenge that remains is how to ensure that the limited input data contains enough scene information, meaning all viewpoints are optimal, which is a critical problem that needs to be addressed.
II-C Uncertainty Estimation in NeRFs
NBV problem[18] aims to select the next viewpoint that maximizes information gain, enhancing scene understanding and minimizing uncertainty. Estimating uncertainty is key for NBV selection, but this remains a challenge in deep learning. Traditional methods like Bayesian Neural Networks (BNN)[27] require significant computational resources, limiting its use in extensive neural architectures. Deep Ensembles[28] need training multiple models and combining their predictions to estimate uncertainty, while MC-Dropout[29] relies on multiple forward passes during testing, leading to inefficiency.
Emerging works address NBV selection for NeRF training under constrained budgets. Some approaches modify NeRF’s architecture, rendering, or training mechanisms, such as S-NeRF[30] and CF-NeRF[19], which use Bayesian methods to model radiance variance and density as random variables, estimating posterior distributions through variational inference. Active-NeRF[20] models radiance as a Gaussian distribution and minimizes negative log-likelihood to predict pixel uncertainty during training. Lee et al.[21] quantify scene uncertainty by calculating the density distribution along sampled rays. Another category of methods does not modify the NeRF architecture itself, making them applicable to any NeRF model. Niko et al.[23] estimate uncertainty via variance of ensemble predictions based on termination probability of rays intersecting unknown areas, while Kopanas et al.[31] introduce observation frequency and angular uniformity to assess camera placement uncertainty. Both methods are based on Instant-NGP. Our innovative approach combines rendering uncertainty in RGB space with spatial uncertainty of camera placement, proposing a hybrid uncertainty estimation for more effective NBV selection.
@@ II-C Uncertainty Estimation in NeRFs
Figure 3: Overview of Our Framework. It consists of three parts: I. Based on the 5D information of the images, we input them into NeRF for training. We define the color at a certain as a gaussian distribution and quantitatively obtain the mean and variance of color from the MLP, generating our rendering uncertainty. II. We extract the positional information of the images from the training set and input them into a classifier to determine their respective Voronoi positional uncertainty, resulting in hybrid uncertainty. III. Based on the hybrid information uncertainty, we select the images with the maximum gain from the candidate set and add them to the training set, iterating continuously.
@@ III Background
NeRF represents a scene as a continuous function $F_{\theta}$ that outputs both the emitted radiance value and volume density. Specifically, given a 3D $\mathbf{x}=(x,y,z)$ in the scene and a viewing direction vector $\mathbf{d}=(\theta,\phi)$, a multi-layer perceptron model is used to generate the corresponding volume density $\sigma$ and color $\mathbf{c}=(r,g,b)$ as follows:
@@
where $\gamma_{x}(\cdot)$ and $\gamma_{d}(\cdot)$ are the positional encoding functions, and $f$ represents the intermediate feature that is independent of the viewing direction $\mathbf{d}$. An interesting observation is that the radiance color is influenced solely by its own 3D coordinates and the viewing direction, rendering it independent of other locations.
@@
$\hat{C}(\mathbf{r})$ $=\sum_{i=1}^{N_{s}}\alpha_{i}c(\mathbf{r}(t_{i})),$ (4)
$\alpha_{i}$ $=\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right)\left(1-\exp(-\sigma_{i}\delta_{i})\right),$ (5)
where $\delta_{i}=t_{i+1}-t_{i}$ is the distance between adjacent samples, and $N_{s}$ denotes the number of samples.Based on this approach, NeRF optimizes the continuous function $F_{\theta}$ by minimizing the squared reconstruction errors between the ground truth RGB images $\{I_{i=1}^{N}\}$ and the rendered pixel colors.
@@
where $\mathbf{r_{i}}$ is a sampled ray, and $C(\mathbf{r_{i}}),\hat{C}^{c}(\mathbf{r_{i}}),\hat{C}^{f}(\mathbf{r_{i}})$ correspond to the ground truth, coarse model prediction, and fine model prediction respectively.
@@ IV IOVS4NeRF
In this paper, we focus on optimal view selection for large-scale NeRFs. In this section, we present the IOVS4NeRF framework in section. IV-A followed by detailed descriptions of the proposed hybrid uncertainty estimation (section. IV-B) and implementation (section. IV-C).
@@ IV-A Framework
Initialization. As shown in Fig. 2, we randomly select a fixed proportion of images (we experimentally recommend 15%) from the input dataset to perform NeRF training for initialization.
Uncertainty Estimation. Note that the IOVS4NeRF selects the best view based a novel information gain metric, named hybrid uncertainty, with two main terms: 1) rendering uncertainty and 2) positional uncertainty:
  • For rendering uncertainty, we input 5D coordinates (x,d) of the remaining photo set into a modified NeRF network with threshold sampling, which then outputs both color c and volume density $\sigma$, then it goes to opacity $\alpha$ and the variance. Subsequently, we calculate the rendering uncertainty of the image by integrating both c and variance into a gaussian distribution.
  • For positional uncertainty, we input 3D coordinates x of the remaining photo set, and a classifier determines whether the trajectory is planar or non-planar. We then estimate the information using Voronoi diagrams to obtain the positional uncertainty.
  • Then, we normalize and sum the two uncertainties to calculate the hybrid uncertainty, as shown in the following formula:
    ${\psi^{2}}(I)=Norm(\sigma_{pos}^{2}(I))+Norm(\sigma_{rgb}^{2}(I)),$ (7)
    where I is the ground truth image while $Norm(\cdot)$ is normalization function. $\sigma_{rgb}^{2}(I)$ and $\sigma_{pos}^{2}(I)$ are rendering and positional uncertainty respectively that defined in Eq. (12) and Eq. (18).
Incremental Selection and Optimization. We select the image with the highest hybrid uncertainty score and continually add it to the training set. This process is repeated until a specific reconstruction effect is achieved or a preset limit on the number of selected images is reached.
@@
IV-B Hybrid Uncertainty Estimation
The proposed hybrid uncertainty (Eq. (7)) includes two terms, namely rendering uncertainty $\sigma_{rgb}^{2}(I)$ and positional uncertainty $\sigma_{pos}^{2}(I)$. Note that our hybrid uncertainty is an epistemic uncertainty [27], which can be reduced by collecting more data.
Rendering Uncertainty. In scenarios with limited training data, the original NeRF framework often results in fragmented outputs due to predicting zero volume density for unobserved regions. To address this issue, we propose representing the RGB values at each scene as a gaussian distribution[20], rather than a single value. The predicted variance indicates the uncertainty at specific locations, allowing the model to assign larger variances to unobserved areas, thereby enhancing information gain for completing renderings and reducing epistemic uncertainty.
Specifically, we define the color at a certain as a gaussian distribution with parameters representing the mean $c(\textbf{r}({t_{i}}))$ and variance ${{\beta}^{2}}(\textbf{r}(t_{i}))$. We establish a variance model by adding an additional branch outside NeRF’s MLP network using information entropy. The information entropy of each sampling point can be considered the variance of the color distribution at that point. This is because higher information entropy indicates more dispersed information distribution. It requires more information to describe, suggesting a greater difference between the predicted result and the true value, which aligns with the definition of variance:
${\beta^{2}}({\bf{r}}({t_{i}}))=-P({\alpha_{i}})\log P({\alpha_{i}}),$ (8)
where ${\alpha_{i}}$ is the same as Eq. (5), and $P(\cdot)$ represents proportion of ${\alpha_{i}}$in the $\sum{{\alpha_{i}}}$ along the ray.
During the rendering process, similar volume rendering methods can be used to handle new neural radiation fields with uncertainty. That is because volume rendering can be seen as a linear combination of sampled points along rays. And the volume density at a specific location is only affected by its 3D coordinates, not by the viewing direction $\mathbf{d}$, which makes the distributions at different s independent of each other. Based on these conditions, if we represent the gaussian distribution at the $\textbf{r}(t_{i})$ then, according to the conjugacy of the gaussian distribution, the rendering values along this ray naturally follow a gaussian distribution as:
$\hat{C}({\textbf{r}})\sim Beta(\bar{C}(\textbf{r}),{{\bar{\beta}}^{2}}(\textbf{r})),$ (9)
where,
$\bar{C}({\textbf{r}})=\sum\limits_{i=1}^{{N_{s}}}{{\alpha_{i}}c({\textbf{r}}({t_{i}})),}$ (10)
${\bar{\beta}^{2}}({\textbf{r}})=\sum\limits_{i=1}^{{N_{s}}}{\alpha_{i}^{2}{\beta^{2}}({\textbf{r}}({t_{i}}))}),$ (11)